What are the x- and y-intercepts for the following equation: 2x+3y=6. Tell how to find each intercept.
Answer by venugopalramana(3286)

2x+3y=6 being a linear equation in x and y represents the graph of a straight line . we have to find its x and y intercepts ,that is where does the above graph cuts the X axis and Y axis.We can do it analytically by noting the fact that the graph intersects the X axis at a point where the point's Y coordinate is zero and ,similarly ,the graph intersects the Y axis at a point where the point's X coordinate is zero.Hence we get the X intercept by putting y=0 in the given equation and solving for x.Similarly we get the Y intercept by putting x=0 in the given equation and solving for y.
Substituting y=0 , we get
2x+3*0=6
2x=6
x=3 is the x intercept.
Substituting x=0 , we get
2*0+3y = 6
3y=6
y=2 is the y intercept.
.